OpenStudy (emilyjones284):

MEDAL AND FAN! (: Given a mean of 8 and a standard deviation of 0.7, what is the z-score of the value 9 rounded to the nearest tenth?

jimthompson5910 (jim_thompson5910):

z = (x - mu)/sigma z = (9 - 8)/0.7 z = ?? note: mu = mean, sigma = standard deviation

OpenStudy (emilyjones284):

1.4?

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (emilyjones284):

thank you!! (:

jimthompson5910 (jim_thompson5910):

np